OpenStudy (blackbird02):

Proving trigonometric identities. Prove that: ((sin(x)-cos(x)-1)/(sin(x)+cos(x)-1))=((cos(x)+1)/(sin(x))

OpenStudy (blackbird02):

\[\frac{ \sin(x)-\cos(x)-1 }{ \sin(x)+\cos(x)-1 }=\frac{ \cos(x)+1 }{ \sin(x) }\]

Miracrown (miracrown):

I suggest we cross multiply then simplify both sides until we get to a true identity.

OpenStudy (calculusfunctions):

Begin with the more complicated side and simplify until you obtain the expression on the other side. Which side do you think you should start with @blackbird02 ?

OpenStudy (blackbird02):

i think we should start with the left side.

OpenStudy (calculusfunctions):

Cross multiplication is a cheap way out for a lack of a better phrase.

Miracrown (miracrown):

Yep. Yep.

OpenStudy (calculusfunctions):

Yes! @blackbird02 start with the left side.

Miracrown (miracrown):

This will not go well if you only work on one side, we work on both sides simultaneously until we get to a true identity.

OpenStudy (blackbird02):

should i do this? \[\frac{ \frac{ 1 }{ \csc(x) }-\frac{ 1 }{ \sec(x) }-1 }{ \frac{ 1 }{ \csc(x) }+\frac{ 1 }{ \sec(x) }-1}=\frac{ \cos(x)-1 }{ \sin(x) }\]

Miracrown (miracrown):

no, it will not help

OpenStudy (calculusfunctions):

Not necessary! No

Miracrown (miracrown):

I was suggesting you to cross multiply... @blackbird02

OpenStudy (blackbird02):

How do I work on the left side? What should I do first?

OpenStudy (blackbird02):

@Miracrown Besides cross multiplying, what could be other steps? We're actually not allowed to cross multiply.

Miracrown (miracrown):

I was saying, we cannot work only on one side we need to work on BOTH sides at the same time.... cross multiply, then simplify.

Miracrown (miracrown):

whatever you do, you will do a form of cross multiplicaiton for this one

OpenStudy (calculusfunctions):

Okay well I see you prefer @Miracrown to help you. Good luck. Bye.

Miracrown (miracrown):

Hey @calculusfunctions, don't be too hasty. =]

OpenStudy (blackbird02):

@calculusfunctions Well, I could use a little advice from you too. What do you think?

Miracrown (miracrown):

I do not think there is any other way except cross multiplication. if you are not allowed to do that, I do not think I am able to solve it. Adios!

OpenStudy (blackbird02):

Other ways I can think is to transform them to other functions like what I posted above, but I'm stuck from there.

OpenStudy (calculusfunctions):

@Miracrown I wasn't being hasty. I just didn't want to waste my time if you're already helping.

Miracrown (miracrown):

Are you tricking us @blackbird02? This is not an identity that is why nothing works, the left side is NOT equal to the right side.

OpenStudy (calculusfunctions):

First @blackbird02 I think you copied the problem incorrectly. The right side should be\[-\frac{ \cos x +1 }{ \sin x }\]

OpenStudy (michele_laino):

@blackbird02 Sorry, may be the right side equal to this? \[-\frac{ 1+\cos(x) }{ \sin(x) }\]

OpenStudy (calculusfunctions):

Right @blackbird02 that's what I just said. Now it works.

OpenStudy (blackbird02):

@calculusfunctions yes, that's right.

OpenStudy (calculusfunctions):

@blackbird02 I'll get you started. First multiply the numerator and the denominator of the left side by sinx + cosx +1. Go ahead and post the result of this multiplication and then I'll walk you through the rest

OpenStudy (calculusfunctions):

Make sure you multiply both the numerator and the denominator by\[\sin x +\cos x +1\]

OpenStudy (blackbird02):

Ok, just a sec

OpenStudy (calculusfunctions):

@Michele_Laino I hope you're not posting a solution! He needs to learn NOT look at someone else's work. Thank you!

Miracrown (miracrown):

Good advice - @calculusfunctions.

OpenStudy (michele_laino):

@calculusfunctions Ok!, done!

OpenStudy (blackbird02):

\[\frac{ \sin^2(x)-\cos^2(x)-2\cos(x)-1 }{ \sin^2(x)+2\sin(x)\cos(x)+\cos^2(x)-1 }\]

OpenStudy (blackbird02):

is this correct?

OpenStudy (calculusfunctions):

Excellent @blackbird02 Now the next step??

OpenStudy (calculusfunctions):

Do recognize the Pythagorean Identity in the denominator @blackbird02 ?

OpenStudy (blackbird02):

It's sin^2A+cos^2B = 1, right?

OpenStudy (calculusfunctions):

Yes!

OpenStudy (calculusfunctions):

Also write\[\sin ^{2}x\]in the numerator in terms of\[\cos ^{2}x\]using the same Pythagorean Identity of course. Go Ahead!!

OpenStudy (calculusfunctions):

@blackbird02 ??

OpenStudy (blackbird02):

@calculusfunctions I'm working on it

OpenStudy (calculusfunctions):

Okay no problem! Just wanted to make sure you're still there. Take your time.

OpenStudy (blackbird02):

I'm not really sure if I did this correctly. Can you check it? \[\frac{ -2\cos^2(x) }{ 2\sin^2(x)+2\sin(x)\cos(x) }\]

OpenStudy (calculusfunctions):

No! From your first step which was correct, you had\[\frac{ \sin ^{2}x -\cos ^{2}x -2\cos x -1}{ \sin ^{2}x +2\sin x \cos x +\cos ^{2}x -1 }\]If you replace\[\sin ^{2}x +\cos ^{2}x =1\]in the denominator, then what should that denominator be now?

OpenStudy (blackbird02):

The denominator would now be \[2\sin(x)\cos(x)\] ?

OpenStudy (calculusfunctions):

Yes!

OpenStudy (calculusfunctions):

And if you write \[\sin ^{2}x\]in the numerator in terms of cosine, then what should the numerator be?

OpenStudy (blackbird02):

The whole fraction would be: \[\frac{ 1-\cos^2(x)-\cos^2(x)-2\cos(x)-1 }{ 2\sin(x)\cos(x) }\]

OpenStudy (calculusfunctions):

Very Good @blackbird02 !! Now collect like terms in the numerator.

OpenStudy (blackbird02):

Simplifying it would be: \[\frac{ -2\cos^2(x)-2\cos(x) }{ 2\sin(x)\cos(x) }\]

OpenStudy (blackbird02):

What do I do next?

OpenStudy (calculusfunctions):

Excellent! Now do you recognize greatest common factor in the numerator?

OpenStudy (blackbird02):

cos(x) ? I'm not really sure.

OpenStudy (calculusfunctions):

What is the greatest common factor in the numerator? Do you not see that -2cosx is common to both terms in the numerator?

OpenStudy (blackbird02):

So, I'll factor it out?

OpenStudy (calculusfunctions):

Yes! Go ahead!

OpenStudy (blackbird02):

Ok, I got it now! @calculusfunctions thank you so much for helping me out and guiding me throughout the process. I really appreciate it.

OpenStudy (calculusfunctions):

No Problem!! Good Luck with everything!